OpenStudy (broner33):

Find the length of the radius of the following circle. (x + 2)2 + y2 = 10 a.5 b.√(10) c.100

OpenStudy (isaidavila):

=r^2 10=r^2 r=√10

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @broner33 @~ You might not have noticed but you are in the "OpenStudy Feedback" subject. Hover your mouse where it says OpenStudy Feedback and find the correct subject by clicking the 'Find More Subjects' button or simply clicking the correct subject in the list that you have already studied. You can also get to the correct subject easily by going to this link: http://www.openstudy.com/study#/groups/Mathematics

OpenStudy (broner33):

thanks:)